Write down the restriction(s) and solve for x: 5/6-3/2x=3/x-2/3
Answer by CubeyThePenguin(3113)

I have answered this already (question #1093997)

5/6 - 3/2x = 3/x - 2/3 (the only restriction is that x can't be equal to 0)
Multiply both sides by 6x.
5x - 9 = 18 - 4x
9x = 27
x = 3